# TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

### v2

@@ Introduction
Cross-Modal Alignment. This module aims to achieve cross-modal alignment between soft-hard prompts and temporal data. Its core design resembles time series reprogramming in semantic space embedding. Specifically, cross-modal alignment is performed separately for soft/hard prompts and time series: (a) Dynamically align adaptive soft prompts with time series via a multi-head self-attention module, obtaining fused representation $Z_{1}$; (b) Process hard prompt text and time series using an identical architecture, obtaining fused representation $Z_{2}$. Finally, the joint representation $Z=[\tilde{X},Z_{1},Z_{2}]$ is formed through element-wise concatenation.
Time Series Forecasting
As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM. Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$. GPT-2 is used by default unless otherwise specified (Radford et al. 2019).
Experiments
To validate the effectiveness of TSF-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting. Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE). All experiments are conducted on V100 GPUs. Additional details (e.g., datasets, SOTA methods, full results) are provided in the appendix.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al. 2023a), TimeCMA (Liu et al. 2024a); (2) Transformer-based: Autoformer (Wu et al. 2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al. 2023), PatchTST (Nie et al. 2022); (3) CNN-based: TimeNet (Wu et al. 2023); (4) RNN-based: Mamba(Gu and Dao 2023); (5) MLP-based: DLinear (Zeng et al. 2023).
Long-term Forecasting
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that TSF-Prompt outperforms all baselines in most cases. LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations. Compared to other LLMs-based approaches, we observe that TSF-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA. This is attributed to our integration of both soft/hard prompts and cross-modal alignment in TSF-Prompt.
@@
ETTh1 0.491 0.469 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
ETTh2 0.388 0.408 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
ETTm1 0.418 0.411 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
ETTm2 0.289 0.333 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
Exchange 0.366 0.410 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
Weather 0.270 0.291 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
munichCarbon $0.419$ $0.472$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
chinaCarbon 0.720 0.610 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
worldCarbon 0.803 0.666 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
Table 4: Zero-shot learning on ETT datasets. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$h1\rightarrow m1$ 96 0.723 0.553 $0.746$ 0.552 $0.772$ $0.565$ $0.766$ $0.573$ 0.713 $0.563$ $0.825$ $0.599$ $0.783$ $0.602$ $0.826$ $0.578$ $0.856$ $0.596$
192 $0.762$ 0.564 $0.832$ $0.587$ $0.807$ $0.581$ 0.749 0.568 0.737 $0.585$ $0.797$ $0.598$ $0.773$ $0.606$ $0.846$ $0.593$ $0.845$ $0.597$
336 0.755 0.557 $0.793$ 0.583 $0.807$ $0.586$ $0.843$ $0.589$ 0.757 $0.604$ $0.798$ $0.610$ $0.786$ $0.617$ $0.835$ $0.596$ $0.829$ $0.594$
720 0.765 0.584 $0.813$ 0.601 $0.819$ $0.601$ $0.865$ $0.605$ 0.800 $0.631$ $0.846$ $0.641$ $0.85$ $0.654$ $0.84$ $0.605$ $0.816$ $0.606$
Avg 0.751 0.564 $0.796$ 0.581 $0.801$ $0.583$ $0.806$ $0.584$ 0.752 $0.596$ $0.816$ $0.612$ $0.798$ $0.620$ $0.837$ $0.593$ $0.836$ $0.598$
$h2\rightarrow m2$ 96 0.213 0.301 0.221 $0.306$ $0.225$ $0.309$ $0.226$ $0.312$ $0.228$ $0.331$ $0.244$ $0.331$ $0.495$ $0.496$ $0.228$ $0.313$ 0.213 0.300
192 0.276 0.337 $0.279$ $0.337$ $0.286$ $0.341$ $0.287$ $0.343$ $0.300$ $0.379$ $0.298$ $0.358$ $0.638$ $0.562$ $0.287$ $0.344$ 0.273 0.334
336 0.330 0.365 0.335 0.370 $0.341$ $0.372$ $0.344$ $0.374$ $0.368$ $0.424$ $0.362$ $0.397$ $0.742$ $0.611$ $0.345$ $0.376$ $0.339$ $0.372$
720 0.432 0.420 $0.433$ 0.422 $0.438$ $0.426$ $0.443$ $0.431$ $0.514$ $0.507$ $0.470$ $0.456$ $1.017$ $0.741$ $0.442$ $0.427$ 0.429 0.420
Avg 0.313 0.356 $0.317$ $0.359$ $0.322$ $0.362$ $0.325$ $0.365$ $0.352$ $0.410$ $0.343$ $0.385$ $0.723$ $0.602$ $0.325$ $0.365$ 0.315 0.356
@@
Results. The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. TSF-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
Zero-shot Forecasting
Setups. LLMs possess the potential for zero-shot time series forecasting. We further investigate the cross-domain zero-shot learning capabilities of LLMs. Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples. Most experimental settings follow the long-term forecasting configuration. We evaluate zero-shot learning performance on ETT-series and carbon emission datasets. Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
Results. Results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting. In the $ETTh1\rightarrow ETTm1$ setting, TSF-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global). Results demonstrate that TSF-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
Table 5: Zero-shot learning on carbon emission datasets. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE
$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
Table 6: Ablations on different components. Bold: the best.
Methods TSF-Prompt w/o Soft Prompt w/o Hard Prompt w/o Alignment
Metric MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
Table 7: Ablations on different hard prompt configurations.
Methods TSF-Prompt w/o Historical Data w/o Statistical Data
Metric MSE MAE MSE MAE MSE MAE
ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
Figure 4: Ablations on different soft prompt parameters.
Figure 5: Ablations on different projection dimensions
Ablation Study
We analyze the effectiveness of dual-path prompting and cross-modal alignment components. As shown in Table 6, ablations are conducted on multiple datasets (ETTm1, ETTh1, munichCarbon, chinaCarbon, worldCarbon). Removing either soft prompts, hard prompts, or cross-modal alignment degrades model performance. Prompt engineering is critical for large models to comprehend inputs: the absence of hard or soft prompts reduces performance by approximately 3.2%. The most severe impact occurs without cross-modal alignment, preventing the LLM from effectively understanding time series and causing a 3.7% performance drop.
Ablations on Hard Prompt. We further investigate how different types of information within hard prompts affect TSF-Prompt’s performance. As shown in Figure 3, we compare TSF-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration. As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃% performance degradation.
Ablations on Soft Prompt. Additionally, we investigate the impact of soft prompt pool size and soft prompt length on TSF-Prompt’s performance. As shown in Figure 4, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon. This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
Ablations on Parameter Sensitivity. We analyze the impact of intermediate projection dimension $d_{model}$ and multi-head attention head dimension $d_{ff}$ on model performance. As illustrated in Figure 4, varying parameter configurations have minimal impact on the ETTh1 dataset, possibly due to its sufficient sample size enabling effective model fitting. In contrast, for the chinaCarbon dataset, specific optimal configurations exist for both $d_{model}$ and $d_{ff}$, enabling optimal information flow within the network.

### v3

@@ Introduction
Cross-Modal Alignment. This module is designed to achieve cross-modal fusion among the time series data, soft prompts, and hard prompts to eliminate the modality gap. Many existing works simply concatenate these variables, which can lead to severe feature redundancy when the number of prompt tokens is large (Jin et al. 2024). Our specific approach is as follows: (a) We use a multi-head cross-attention mechanism to align the soft prompt with the time series, obtaining a fused representation $Z_{1}$. Similarly, the hard prompt is aligned with the time series to obtain $Z_{2}$. (b) We then fuse the original time series and the fused representations $Z_{1}$ and $Z_{2}$ via a gating mechanism, yielding the final representation $Z$.
$Z=g_{0}\odot\tilde{X}+g_{1}\odot Z_{1}+g_{2}\odot Z_{2}$ (5)
@@
ETTh1 0.464 0.454 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
ETTh2 0.387 0.406 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
ETTm1 0.400 0.404 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
ETTm2 0.285 0.329 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
Exchange 0.358 0.403 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
Weather 0.268 0.288 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
munichCarbon $0.416$ $0.469$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
chinaCarbon 0.654 0.577 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
worldCarbon 0.738 0.619 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
Table 4: Ablation Experiment
Methods Time-Prompt w/o SP w/o HP w/o CMA w/o LoRA Qwen 3 0.6B
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.430 0.439 $0.434$ $0.441$ $0.437$ $0.442$ $0.434$ $0.442$ $0.435$ $0.440$ $0.717$ $0.579$
ETTm1 0.380 0.398 $0.383$ $0.399$ $0.384$ $0.400$ $0.386$ $0.402$ $0.386$ $0.401$ $0.719$ $0.564$
munichCarbon 0.411 0.532 $0.413$ $0.466$ $0.414$ $0.467$ $0.414$ $0.467$ $0.416$ $0.468$ $0.652$ $0.588$
chinaCarbon 0.571 0.532 $0.581$ $0.540$ $0.584$ $0.540$ $0.587$ $0.544$ $0.583$ $0.538$ $0.802$ $0.652$
worldCarbon 0.617 0.544 $0.636$ $0.558$ $0.633$ $0.556$ $0.633$ $0.556$ $0.630$ $0.552$ $0.869$ $0.713$
Time Series Forecasting
Many existing works rely solely on frozen LLMs. Given the high dimensionality of the LLM’s semantic space, they often resort to direct truncation to prevent the curse of dimensionality. This approach, however, is detrimental to activating the LLM’s potential for time series forecasting. In this module, we first employ LoRA to efficiently fine-tune the LLM, enhancing its ability to comprehend continuous values. Second, to address the excessively high semantic dimensionality, we project its representations into a lower-dimensional space, extract global features, and then jointly project these global features together with local features to produce the final prediction $Y$.
$F_{\text{local}}$ $=\text{Linear}(Z),$ (6)
$F_{\text{global}}$ $=\text{Mean}(F_{\text{local}}),$ (7)
$F_{\text{fused}}$ $=\text{Linear}([F_{\text{local}};F_{\text{global}}]),$ (8)
$Y$ $=\text{Head}(\text{Mean}(F_{\text{fused}})).$ (9)
Experiments
Dataset. We utilize six widely-used public datasets covering multiple domains, including electricity, finance, and weather, as well as three carbon emission datasets with different spatial scales. Among them, the Munich carbon emission dataset was collected by one of our co-authors. We conduct extensive experiments across these nine datasets to demonstrate the superiority of Time-Prompt, including tasks such as long-term forecasting and few-shot forecasting. Zero-shot forecasting is not considered in this work, as we think LLM-based frameworks currently lack strong generalization capability in such settings.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al. 2024), TimeCMA (Liu et al. 2025a); (2) Transformer-based: Autoformer (Wu et al. 2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al. 2023), PatchTST (Nie et al. 2022); (3) CNN-based: TimeNet (Wu et al. 2023); (4) RNN-based: Mamba(Gu and Dao 2023); (5) MLP-based: DLinear (Zeng et al. 2023).
Implementation. Since the vast majority of LLM-based time series forecasting methods adopt GPT-2, we also use GPT-2 by default in this work for fair comparison (Radford et al. 2019). The evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE). Additional details (e.g., datasets, baseline, full results) are provided in the appendix.
Long-term Forecasting
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that Time-Prompt outperforms all baselines in most cases. Time-Prompt achieves the best performance in 14 out of 18 benchmark settings. Overall, LLM-based methods demonstrate superior performance compared to conventional deep learning approaches. Specifically, Time-Prompt reduces MSE and MAE by 5.5% and 2.7% over Time-LLM, by 6.3% and 4.2% over TimeCMA, and by 4.3% and 2.4% compared to PatchTST—the strongest deep learning baseline.
@@
Results. The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. The superiority of Time-Prompt is even more pronounced in few-shot scenarios compared to long-term forecasting. Time-Prompt achieves the best performance in 30 out of 36 benchmark settings Under the 10% training data setting, Time-Prompt reduces MSE and MAE by 7.1% and 4.3% over Time-LLM, and by 8.0% and 5.5% over TimeCMA. Under the more challenging 5% training data setting, it further achieves 8.8% and 5.7% MSE/MAE reductions compared to Time-LLM, and 7.9% and 5.5% reductions compared to TimeCMA.
Table 5: Modal Analysis.
soft prompt 3 5 10 15
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.376 0.402 0.376 0.402 0.374 0.401 0.377 0.402
munichCarbon 0.238 0.352 0.239 0.352 0.237 0.351 0.236 0.351
chinaCarbon 0.473 0.472 0.462 0.469 0.458 0.469 0.467 0.473
hard prompt 3 5 10 15
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.378 0.405 0.376 0.402 0.376 0.402 0.375 0.400
munichCarbon 0.236 0.350 0.239 0.352 0.242 0.354 0.236 0.350
chinaCarbon 0.464 0.474 0.462 0.469 0.466 0.469 0.460 0.467
d_model 32 64 128 256
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.376 0.402 0.378 0.402 0.378 0.402 0.381 0.403
munichCarbon 0.239 0.352 0.236 0.349 0.241 0.352 0.238 0.350
chinaCarbon 0.462 0.469 0.467 0.473 0.470 0.477 0.455 0.468
d_ff 32 64 128 256
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.381 0.405 0.380 0.406 0.376 0.402 0.380 0.400
munichCarbon 0.239 0.352 0.243 0.354 0.239 0.352 0.239 0.351
chinaCarbon 0.482 0.483 0.460 0.469 0.462 0.469 0.464 0.471
Ablation Experiment
As shown in Table 4, we conduct an ablation study on multiple datasets (ETTh1, ETTm1, munichCarbon, chinaCarbon, worldCarbon) to analyze the effects of soft prompts (SP), hard prompts (HP), cross-modal alignment (CMA), and LoRA. All experiments follow the long-term forecasting setting, using four different prediction horizons, and the reported results are averaged across these settings. The results indicate that all these modules activate the LLM’s time series forecasting capabilities to varying degrees. Compared to the prompting engineering, LoRA and cross-modal alignment have a more significant impact on the LLM’s forecasting ability. Additionally, we replace GPT-2 with the Qwen-3 0.6B model, which lead to a significant performance drop. We hypothesize that this is because larger-parameter models struggle to effectively handle continuous-valued time-series data. This observation aligns with findings in the time-series foundation model literature: most published time-series foundation models are relatively small, typically in the few-hundred-million parameter range or even smaller (Team 2025). This also indicates that in future research, we should focus more on the design of model architecture and the tuning of model parameters, rather than solely on prompt engineering. Furthermore, greater attention should be paid to foundation models and small-scale LLMs (Liu et al. 2025b; Shi et al. 2025).
Modal Analysis
We experiment with the settings for soft prompt length, hard prompt length, $d_{model}$, and $d_{ff}$ on multiple datasets (ETTh1, munichCarbon, chinaCarbon). The hard prompt length refers to the number of tokens retained after the hard prompt is encoded by the LLM. Due to the relatively large number of tokens, we truncate it to mitigate the curse of dimensionality. This approach is supported by existing literature, which indicates that the key information is predominantly located in the latter part (Liu et al. 2025a). Here, $d_{model}$ represents the intermediate dimension of the patch reprogram module, while $d_{ff}$ is the intermediate dimension of the projection module.
As shown in the Table 5, the soft prompt length should neither be too short nor too long. Since soft prompts are designed to guide the LLM’s behavior, an excessively long sequence of tokens is unnecessary. In contrast, hard prompts are constructed by textualizing time-series data, so retaining more tokens is generally better, as truncation inevitably leads to information loss. However, considering the trade-off between efficiency and performance, we opt to keep only the most informative segment. For $d_{ff}$ and $d_{model}$, we employ a grid search to select the optimal parameter combination. Ultimately, for all datasets, we default to $d_{ff}=128$ and $d_{model}=32$.